### v2

@@ Abstract
In spring 1949 about 70 physicists from eight countries met in Florence to discuss recent trends in statistical mechanics. This scientific gathering, co-organized by the Commission on Thermodynamics and Statistical Mechanics of the International Union of Pure and Applied Physics (IUPAP) and the Italian Physical Society (SIF), initiated a tradition of IUPAP-sponsored international conferences on statistical mechanics that lasts to this day. In 1977, when this conference series took the name of StatPhys, the foundational role of the Florence conference was recognized by retrospectively naming it StatPhys1. This paper examines the dual scientific and social significance of the conference, situating it in the broader contexts of the post-World War II reconstruction in Italian physics and of the revitalization of the international science organization. Through an analysis of IUPAP archives and Italian records, we illustrate how the event’s success hinged on the aligned objectives of its organizers. Internationally, it was instrumental in defining the scientific and organizational foundations for the activities of IUPAP commissions during a critical phase of IUPAP’s history, when the Union was resurging on the international stage post-interwar period inactivity. Nationally, the conference served as a cornerstone in SIF’s strategy to re-establish Italian physics’ international stature and to aid the domestic revitalization of physics through the internationalization of its activities, notably of its flagship journal, Il Nuovo Cimento. This analysis not only sheds light on the conference’s impact but also informs recent discussions in the history of science about the multiple roles of international scientific conferences.
@@ 1 Introduction
On May 17th 1949, about 70 physicists [^1] Almost all males, how often occurred in that historical period for these sorts of international scientific gatherings. from the United States and Europe, almost uniquely Western, met at Villa Favard in Florence for the opening session of a conference devoted to discuss frontier research topics from the statistical mechanics of interacting systems (see Figs. 1 and 2). [^2] Giovanni Polvani to Mons. Arnoux, March 24, 1949, Box 3, Folder 2, Presidenza Polvani, Archives of the Italian Physical Society (hereafter ASIF). This 4-day conference initiated a regular series of increasingly larger international meetings dedicated to statistical physics that lasts to this day: the StatPhys international conferences sponsored by the International Union of Pure and Applied Physics (IUPAP). In the 1970s, when this tradition was well established and StatPhys emerged as the official name of these regular conferences, the role of the Florence conference was ratified by considering it StatPhys1.333In the documents of the IUPAP archives the first mention of ‘StatPhys’ occurs in relation to the conference held in Haifa in 1977, called StatPhys13; see, e.g., IUPAP General Report 1979 p. 54, Series B2aa, Vol. 2, IUPAP Archives, Gothenburg Secretariat (hereafter IuG), Center for the History of Science, Royal Swedish Academy of Science; see also the proceedings of the conference [4]. However, we do not exclude that the ‘StatPhys’ label was informally used before the 1977 conference.
@@
As the first of an important disciplinary tradition, it played a role in the international development of an important sub-field of physics, warranting a deeper historical examination than it has received to date. Even more so when one recognizes that the conference was the result of two parallel (re-)construction processes in the post-World War II period: the re-establishment of the international organization of physics and the attempts to (re-)build Italian physics. It was a collaborative effort between the newly revitalized Italian Physical Society (SIF, from its Italian name Società  Italiana di Fisica) and the IUPAP Commission on Thermodynamics and Statistical Mechanics, which was notable for being the first specialized commission ever established by IUPAP, just two years prior. The conference’s pioneering role underscores its significance not only as a platform that fostered the development of an international sub-disciplinary framework within the broader scientific discipline of physics but also as a means for the Italian scientific community to reassert its prominence in the international arena after the war’s devastations.
@@
International scientific conferences have garnered increasing interest from historians of science as key objects for exploring the cultural, social, and political dimensions of the scientific enterprise. A recent European project has shed light on the vast diversity of these gatherings, introducing a taxonomy that categorizes them into four distinct types: disciplinary conferences, which cultivate communities within scientific fields and establish common standards; scientific associations’ conferences primarily aimed at community building; technical conferences, which facilitate interaction between scientists and technicians in inter-governmental contexts to address technical or political challenges; and small ‘elite’ conferences sponsored by wealthy philanthropists, aimed at reinforcing politically-charged research agendas [8]. [^4] See the website’s project at https://heranet.info/projects/public-spaces-culture-and-integration-in-europe/the-scientific-conference-a-social-cultural-and-political-history/. According to the project’s team, despite their varied formats and objectives, these conferences share certain characteristics that render them indispensable to the scientific community. Drawing upon Durkheim’s analysis of religious gatherings, historian Geert Somsen argues that scientific conferences enact rituals and routines that have purely cultural and social goals. They enable participants to celebrate their collective identity and engage with a shared sense of purpose and solidarity [55]. Another important aspect highlighted by the project is that scientific conferences may have geopolitical goals too, as they were, and still are, recognized by larger institutions as valuable tools for fostering international relations [50]. These insights provide a comprehensive framework for examining international scientific gatherings, revealing their multifaceted roles, norms, and practices. This approach offers a lens through which to view specific historical instances, such as the conference under discussion, but it also allows to articulate and complement that framework looking at a different historical case. In fact, the conference presents a special case of “disciplinary conference" type of the above mentioned taxonomy, one mostly defined by its focus on a specific sub-discipline within a larger disciplinary field in relation to an institutional framework devoted to the entire discipline.
@@
The 1949 conference was co-organized by IUPAP, the major international organization devoted to physics. Recently, there has been a growing interest in the study of scientific organizations within the framework of science diplomacy history. [^5] For historiographical analyses of current historiographical trends on science diplomacy, see [58, 5]. The crucial public and military significance of physics immediately following World War II has highlighted IUPAP as a subject of particular interest for analysis through the lens of science diplomacy. Recent historical research has revealed the intricate interactions between scientific endeavors and political strategies that influenced IUPAP’s operations during the early Cold War [34]. These analyses are essential for understanding the broader international context surrounding the Florence conference. Among such studies, Japanese historian of science Kenji Ito has illustrated the essential role international conferences played in the maintenance of knowledge infrastructure and in allowing local scientific communities to rebuild ties after the end of World War II. This was the case of the International Conference of Theoretical Physics held in 1953 Japan, which was pivotal for the local physics community to re-appropriate a central space in the international organization of science that was being rebuilt [25].
@@
To discuss the multiple elements embedded in the organization of the conference, we organize the paper as follows. We start by sketching the prehistory of the event, first by summarizing the development of IUPAP in the interwar period in Section 2, and, in Section 3, by delving into the political issues characterizing Italian changing relations with IUPAP in that period. In Section 4, we contextualize the foundation of the Commission on Thermodynamics and Statistical Mechanics in the post-World War II refoundation of IUPAP and of the international organization of science. In Section 5, we discuss the Florence conference as a key moment in the international (re-)construction of Italian physics after World War II. In Section 6, we analyze the scientific content of the conference and its role in the process of discipline-formation for an international community of statistical physicists. In Section 7, we report some major transformations in the following developments of the Commission and its conferences up to the establishment of the name StatPhys13 for the 1977 conference held in Haifa. In the conclusion (Section 8), we underscore the impact and role of the conference for both the international community of statistical physicists and for the local community of Italian physicists, discussing the implications of our case study for the cultural, social, political and scientific history of scientific conferences.
@@ 2 IUPAP in the interwar period
To accurately contextualize the 1949 Florence conference, it is essential to consider both the general evolution of IUPAP during the interwar period and the specific dynamics of Italian participation, as these aspects were significantly influenced by the shifting political landscapes of the time. IUPAP was funded in 1922 within the International Research Council (IRC), the main framework for the re-organization of international science after World War I. [^6] For the history of IUPAP in the interwar period see [16, 40]. Like all scientific unions of the IRC family, the structure and membership of IUPAP was shaped by the punitive attitude of the victorious nations against the Central Powers. At the time of the IRC founding in 1919, the exclusion of countries belonging to the Central Powers was decreed by statute. A period of twelve years was fixed before this rule could be modified. Only when the original convention expired, in December 1931, would it be possible to amend the statutes and admit the countries initially excluded. These rules applied to all unions that operated within the IRC, as only institutions representing countries that were already members of the IRC could be admitted as members of the unions. This rule was based on the notion of national membership according to which the scientific members of both the IRC and the unions were official national representatives. [^7] The institutional nature of national representatives varied greatly. For instance, a 1928 report states “out of the thirty-five countries which have joined the International Research Council, fourteen are represented by their scientific academies, six by national research councils composed of representatives of the national academies, one by a scientific society, and seven others by a scientific department connected with its government. In seven cases only out of the thirty-five is the government the adhering body.” See [1], p. 390.
@@
At the initial General Assembly of the IRC in 1919, a preliminary statute for a Union of Physics was drafted, but it wasn’t until 1922 that a temporary ten-member Executive Committee was formed to officially establish the Union.888“Union internationale de physique pure et appliquée. Procés-Verbal Assemblée Générale Constitutive, Paris, Décembre 1923,” Series B2aa, Vol. 1, IuG. This early phase of IUPAP’s history was significantly influenced by the IRC’s punitive stance towards the Central Powers, which impacted IUPAP’s operations until the end of World War II. Relative to other scientific unions, such as the International Astronomical Union (IAU), IUPAP was marked by periods of inactivity and diminished influence during the interwar years. The limitations of an organization devoted to international cooperation based on the boycott of German scientists became evident especially in the field of physics, where major developments in theoretical physics had German-speaking scientific communities as their nerve center. Furthermore, the evolving diplomatic landscape, underscored by the Locarno Pact of 1925 between France and Germany and Germany’s subsequent admission into the League of Nations in 1926, highlighted the outdated and scientifically detrimental nature of excluding German scientists from the IRC and its unions, against the backdrop of the political détente of that era.
@@
Although modifications to the IRC statutes in 1926 nominally allowed for the membership of League of Nations countries, potentially paving the way for Germany’s inclusion in the IRC and its unions, the intransigent positions of French and Belgian scientists, on the one side, and of German academies, on the other, prevented this integration [52]. This deadlock had significant repercussions for IUPAP, which lagged in fostering full international cooperation with German physicists, in contrast to the broader trend of re-engagement seen in key meetings like the Como and Solvay conferences of 1927. IUPAP remained characterized by an extremely partial international cooperation despite the fact that William Bragg (president of IUPAP from 1922 to 1931), the French physicist Henri Abraham (IUPAP secretary general from 1922 until his death at Auschwitz in 1944), and other members of IUPAP’s Executive Committee had argued in favor of an early entry of German physicists, even proposing to suspend assemblies until the full participation of German physicists was feasible.999“Union internationale de physique pure et appliquée. Deuxième assemblée générale. Bruxelles, 7 jullet 1925,” Series B2aa, Vol. 1, IuG. As a testament to this stance, IUPAP did not convene any conferences between its second General Assembly in 1925 and 1931, coinciding with the year when the IRC and its unions finally amended their statutes to facilitate broader membership.
@@
When the IRC transformed into the International Council of Scientific Unions (ICSU) in 1931, there was a widespread anticipation that German scientific institutions would swiftly become members of ICSU or its individual unions. However, this integration did not materialize as expected. A significant portion of the German scientific community remained reluctant to join an international body that was originally established with the intention of excluding them. Compounding this issue, disagreements within German academic circles about which institution legitimately represented the German state, along with the evolving economic and political landscape (marked by the aftermath of the 1929 stock market crash and the Nazi party’s ascendancy in 1933) further complicated Germany’s potential membership in ICSU.
@@
IUPAP resumed its activities in 1931 with the hope that German physicists would soon become part of the organization. At the General Assembly that year, U.S. physicist and Nobel Laureate, Robert Millikan, became President of IUPAP with this ambition in mind and grand plans for the extension of the IUPAP activities, including the establishment of the first general scientific commissions, the most relevant of which was the Commission on Symbols, Units and Nomenclature (SUN) aimed to establish international standards. Despite these intentions, the anticipated integration of German physicists failed to materialize, significantly hindering the progress of these initiatives from the beginning. A clear sign of these difficulties was that IUPAP didn’t even have a President from 1934 to 1937. After learning that he had been elected in absentia at the IUPAP 1934 General Assembly as Millikan’s successor, Niels Bohr declined the position, stating that he could not be president of an organization that did not embody the principle of being “truly international." [^10] N. Bohr to H. Abraham, December 12, 1934, Series E1, Vol. 5, Folder 38, IUPAP Archives, Quebec Secretariat (hereafter IuQ), Center for the History of Science, Royal Swedish Academy of Science. After Enrico Fermi also declined, Swedish physicist Manne Siegbahn accepted the position still hopeful of engaging German physicists, but the prevailing political conditions thwarted any significant revival of IUPAP’s endeavors before the onset of World War II brought its activities to a complete standstill.

### v3

@@ Abstract
In spring 1949 about 70 physicists from eight countries met in Florence to discuss recent trends in statistical mechanics. This scientific gathering, co-organized by the Commission on Thermodynamics and Statistical Mechanics of the International Union of Pure and Applied Physics (IUPAP) and the Italian Physical Society (SIF), initiated a tradition of IUPAP-sponsored international conferences on statistical mechanics that lasts to this day. In 1977, when this conference series took the name of StatPhys, the foundational role of the Florence conference was recognized by retrospectively naming it StatPhys1. This paper examines the dual scientific and social significance of the conference, situating it in the broader contexts of the post-World War II reconstruction in Italian physics and of the revitalization of the international science organization. Through an analysis of IUPAP archives and Italian records, we illustrate how the event’s success hinged on the aligned objectives of its organizers. Internationally, it was instrumental in defining the scientific and organizational foundations for the activities of IUPAP commissions during a critical phase of IUPAP’s history, when the Union was resurging on the international scene after the inactivity of the interwar period. Nationally, the conference served as a cornerstone in SIF’s strategy to re-establish Italian physics’ international stature and to aid the domestic revitalization of physics through the internationalization of its activities, notably of its flagship journal, Il Nuovo Cimento. This analysis not only sheds light on the conference’s impact but also informs recent discussions in the history of science about the multiple roles of international scientific conferences.
@@ 1 Introduction
On May 17th 1949, about 70 physicists [^1] Almost all males, how often occurred in that historical period for these sorts of international scientific gatherings. from the United States and Europe, almost uniquely Western, met at Villa Favard in Florence for the opening session of a conference devoted to discuss frontier research topics from the statistical mechanics of interacting systems (see Figs. 1 and 2). [^2] Giovanni Polvani to Mons. Arnoux, March 24, 1949, Box 3, Folder 2, Presidenza Polvani, Archives of the Italian Physical Society (hereafter ASIF). This 4-day conference initiated a regular series of increasingly larger international meetings dedicated to statistical physics that lasts to this day: the StatPhys international conferences sponsored by the International Union of Pure and Applied Physics (IUPAP). In the 1970s, when this tradition was well established and StatPhys emerged as the official name of these regular conferences, the role of the Florence conference was ratified by considering it StatPhys1.333In the documents of the IUPAP archives the first mention of ‘StatPhys’ occurs in relation to the conference held in Haifa in 1977, called StatPhys13; see, e.g., IUPAP General Report 1979, p. 54, Series B2aa, Vol. 2, IUPAP Archives, Gothenburg Secretariat (hereafter IuG), Center for the History of Science, Royal Swedish Academy of Sciences; see also the proceedings of the conference [15]. However, we do not exclude that the ‘StatPhys’ label was informally used before the 1977 conference.
@@
As the first of an important disciplinary tradition, it played a foundational role in the international development of a major sub-field of physics, and thus warrants a deeper historical examination than it has received to date. Even more so when one recognizes that the conference was the result of two parallel re-construction processes in the post-World War II period: the re-establishment of the international organization of physics and the attempts to (re-)build Italian physics. It was a collaborative effort between the newly revitalized Italian Physical Society (SIF, from its Italian name Società Italiana di Fisica) and the IUPAP Commission on Thermodynamics and Statistical Mechanics, which was notable for being the first specialized commission ever established by IUPAP, just two years prior. The conference’s pioneering role underscores its significance not only as a platform that fostered the development of an international sub-disciplinary framework within the broader scientific discipline of physics but also as a means for the Italian scientific community to reassert its prominence in the international arena after the war’s devastations.
@@
International scientific conferences have garnered increasing interest from historians of science as key objects for exploring the cultural, social, and political dimensions of the scientific enterprise. The recent European project The Scientific Conference: A Social, Cultural and Political History has shed light on the vast diversity of these gatherings, introducing a taxonomy that categorizes them into four distinct types: disciplinary conferences, which cultivate communities within scientific fields and establish common standards; scientific associations’ conferences primarily aimed at community building; technical conferences, which facilitate interaction between scientists and technicians in inter-governmental contexts to address technical or political challenges; and small ‘elite’ conferences sponsored by wealthy philanthropists, aimed at reinforcing politically-charged research agendas [6]. [^4] See the website’s project at https://heranet.info/projects/public-spaces-culture-and-integration-in-europe/the-scientific-conference-a-social-cultural-and-political-history/. According to the project’s team, despite their varied formats and objectives, these conferences share certain characteristics that render them indispensable to the scientific community. Drawing upon Émile Durkheim’s analysis of religious gatherings, historian Geert Somsen argues that scientific conferences enact rituals and routines that have purely cultural and social goals. They enable participants to celebrate their collective identity and engage with a shared sense of purpose and solidarity [76]. Another important aspect highlighted by the project is that scientific conferences may have geopolitical goals too, as they were, and still are, recognized by larger institutions as valuable tools for fostering international relations [70]. These insights provide a comprehensive framework for examining international scientific gatherings, revealing their multifaceted roles, norms, and practices. This approach offers a lens through which to view specific historical instances, such as the conference under discussion, but it also allows to articulate and complement that framework looking at a different historical case. In fact, the Florence conference presents a special case of the “disciplinary conference" type of the above mentioned taxonomy, one mostly defined by its focus on a specific sub-discipline within a larger disciplinary field in relation to an institutional framework devoted to the entire discipline.
@@
The 1949 conference was co-organized by IUPAP, the major international organization devoted to physics. Recently, there has been a growing interest in the study of scientific organizations within the framework of science diplomacy history. [^5] For historiographical analyses of current historiographical trends on science diplomacy, see [e.g., 80, 1]. The crucial public and military significance of physics immediately following World War II has highlighted IUPAP as a subject of particular interest for analysis through the lens of science diplomacy. Recent historical research has revealed the intricate interactions between scientific endeavors and political strategies that influenced IUPAP’s operations during the early Cold War [50]. These analyses are essential for understanding the broader international context surrounding the Florence conference. Among such studies, Japanese historian of science Kenji Ito (2024) has illustrated the essential role international conferences played in the maintenance of knowledge infrastructure and in allowing local scientific communities to rebuild ties after the end of World War II. This was the case of the International Conference of Theoretical Physics held in 1953 Japan, which was pivotal for the local physics community to re-appropriate a central space in the international organization of science that was being rebuilt.
@@
To discuss the multiple elements embedded in the organization of the conference, we organize the paper as follows. We start by sketching the prehistory of the event, first by summarizing the development of IUPAP in the interwar period in Section 2, and, in Section 3, by delving into the political issues characterizing Italian changing relations with IUPAP in that period. In Section 4, we contextualize the foundation of the Commission on Thermodynamics and Statistical Mechanics in the post-World War II refoundation of IUPAP and of the international organization of science. In Section 5, we discuss the Florence conference as a key moment in the international (re-)construction of Italian physics after World War II. In Section 6, we analyze the scientific content of the conference and its role in the process of discipline-formation for an international community of statistical physicists. In Section 7, we report some major transformations in the following developments of the Commission and its conferences up to the establishment of the name StatPhys13 for the 1977 conference held in Haifa. In the conclusion (Section 8), we review the impact and role of the conference for both the international community of statistical physicists and for the local community of Italian physicists, discussing the implications of our case study for the cultural, social, political and scientific history of scientific conferences. This paper contributes to filling a historiographical gap in the 20th-century history of physics by exploring a significant episode in the development of statistical mechanics and statistical physics post-World War II, an area that has not been comprehensively investigated yet.
@@ 2 IUPAP in the interwar period
To accurately contextualize the 1949 Florence conference, it is essential to consider both the general evolution of IUPAP during the interwar period and the specific dynamics of Italian participation, as these aspects were significantly influenced by the shifting political landscapes of the time. IUPAP was funded in 1922 within the International Research Council (IRC), the main framework for the re-organization of international science after World War I. [^6] For the history of IUPAP in the interwar period see [24, 57]. Like all scientific unions of the IRC family, the structure and membership of IUPAP was shaped by the punitive attitude of the victorious nations against the Central Powers. At the time of the IRC founding in 1919, the exclusion of countries belonging to the Central Powers was decreed by statute. A period of twelve years was fixed before this rule could be modified. Only when the original convention expired, in December 1931, would it be possible to amend the statutes and admit the countries initially excluded. These rules applied to all unions that operated within the IRC, as only institutions representing countries that were already members of the IRC could be admitted as members of the unions. This rule was based on the notion of national membership according to which the scientific members of both the IRC and the unions were official national representatives. [^7] The institutional nature of national representatives varied greatly. For instance, a 1928 report reads: “out of the thirty-five countries which have joined the International Research Council, fourteen are represented by their scientific academies, six by national research councils composed of representatives of the national academies, one by a scientific society, and seven others by a scientific department connected with its government. In seven cases only out of the thirty-five is the government the adhering body” [38, p. 390].
@@
At the initial General Assembly of the IRC in 1919, a preliminary statute for a Union of Physics was drafted, but it wasn’t until 1922 that a temporary ten-member Executive Committee was formed to officially establish the Union.888“Union internationale de physique pure et appliquée. Procés-Verbal Assemblée Générale Constitutive, Paris, Décembre 1923,” Series B2aa, Vol. 1, IuG. Relative to other scientific unions, such as the International Astronomical Union (IAU), IUPAP was marked by periods of inactivity and diminished influence during the interwar years. The limitations of an organization devoted to international cooperation based on the boycott of German scientists became evident especially in the field of physics, where major developments in theoretical physics had German-speaking scientific communities as their nerve center. Furthermore, the evolving diplomatic landscape, epitomized by the Locarno Pact of 1925 between France and Germany and Germany’s subsequent admission into the League of Nations in 1926, highlighted the outdated and scientifically detrimental nature of excluding German scientists from the IRC and its unions, against the backdrop of the political détente of that era.
@@
Although modifications to the IRC statutes in 1926 nominally allowed for the membership of League of Nations countries, potentially paving the way for Germany’s inclusion in the IRC and its unions, the intransigent positions of French and Belgian scientists, on the one side, and of German academies, on the other, prevented this integration [73]. This deadlock had significant repercussions for IUPAP, which lagged in fostering full international cooperation with German physicists, in contrast to the broader trend of re-engagement seen in key meetings like the Como and Solvay conferences of 1927. IUPAP remained characterized by an extremely partial international cooperation despite the fact that William Bragg (president of IUPAP from 1922 to 1931), the French physicist Henri Abraham (IUPAP secretary general from 1922 until his death at Auschwitz in 1944), and other members of IUPAP’s Executive Committee had argued in favor of an early entry of German physicists, even proposing to suspend assemblies until the full participation of German physicists was feasible.999“Union internationale de physique pure et appliquée. Deuxième assemblée générale. Bruxelles, 7 juillet 1925,” Series B2aa, Vol. 1, IuG. As a testament to this stance, IUPAP did not convene any conferences between its second General Assembly in 1925 and 1931, coinciding with the year when the IRC and its unions finally amended their statutes to facilitate broader membership.
@@
When the IRC transformed into the International Council of Scientific Unions (ICSU) in 1931, there was a widespread anticipation that German scientific institutions would swiftly become members of ICSU or its individual unions. However, this integration did not materialize as expected. A significant portion of the German scientific community remained reluctant to join an international body that was originally established with the intention of excluding them. Compounding this issue, disagreements between German academies about which one legitimately represented the German state, along with the evolving economic and political landscape—marked by the aftermath of the 1929 stock market crash and the Nazi Party’s ascendancy in 1933—further complicated Germany’s potential membership in ICSU.
@@
IUPAP resumed its activities in 1931 with the hope that German physicists would soon become part of the organization. At the General Assembly that year, U.S. physicist and Nobel Laureate, Robert Millikan, became president of IUPAP with this ambition in mind and grand plans for the extension of the IUPAP activities, including the establishment of the first general scientific commissions, the most relevant of which was the Commission on Symbols, Units and Nomenclature (SUN) aimed to establish international standards. When the anticipated integration of German physicists failed to materialize, these initiatives were hindered from the beginning. A clear sign of these difficulties was that IUPAP didn’t even have a president from 1934 to 1937. After learning that he had been elected in absentia at the IUPAP 1934 General Assembly as Millikan’s successor, Niels Bohr declined the position, stating that he could not be president of an organization that did not embody the principle of being “truly international." [^10] N. Bohr to H. Abraham, December 12, 1934, Series E1, Vol. 5, Folder 38, IUPAP Archives, Quebec Secretariat (hereafter IuQ), Center for the History of Science, Royal Swedish Academy of Sciences. In 1937, Enrico Fermi also declined the position without providing any reasons. [^11] Fermi to Abraham, October 4, 1937, Series E1, Vol. 5, Folder 38, IuQ. [^12] It is useful to note that the 1937 offer to Enrico Fermi to become president of IUPAP demonstrates his high international standing at the time, confirmed by the awarding of the Nobel Prize one year later. It also shows that Fermi, like Bohr and Siegbahn, was considered close enough to the German physics community to be able to involve them in IUPAP activities. Only later did Swedish physicist and Nobel Laureate Manne Siegbahn become the new president. At the time, Siegbahn was establishing the first research institute for nuclear physics in Sweden, a country with strong cultural and scientific ties with Germany. He became president still hopeful of engaging German physicists, but the prevailing political conditions thwarted any significant revival of IUPAP’s endeavors before the onset of World War II brought its activities to a complete standstill.